OpenStudy (anonymous):

Select the equations that are parallel and perpendicular to y = one over fourx + 7 and that pass through the point (−8, 9).

OpenStudy (anonymous):

@Falco276 @GParsley @KawasumiKimiko @Luis_Rivera @KawasumiKimiko @Hero @Luis_Rivera

OpenStudy (anonymous):

@brantley28 @Reyespan @saifoo.khan @dan815 @countonme123 @surjithayer

OpenStudy (anonymous):

are there answer choices cause there are many ways to answer it

OpenStudy (anonymous):

\[is the eq. y=\frac{ 1 }{ 4}\left( x+7 \right)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Yes parallel: y = one over fourx − 11 perpendicular: y = −4x + 23 parallel: y = −4x + 11 perpendicular: y = one over fourx − 4 parallel: y = 4x − 7 perpendicular: y = negative one over fourx − 7 parallel: y = one over fourx + 11 perpendicular: y = −4x − 23

OpenStudy (anonymous):

eq of any line parallel to this line is \[y=\frac{ 1 }{ 4 }x+c\] \[becuse \it passes through \left( -8,9 \right)\] 9=-8/4+c c=9+2=11 put the value of c and get the eq.

OpenStudy (anonymous):

so whats the answer

OpenStudy (anonymous):

@surjithayer

OpenStudy (anonymous):

\[y=\frac{ 1 }{ 4 }x+11\]

OpenStudy (anonymous):

slope of the line perpendicular to the given line=-1/m=-4 eq. of any line with soe 4 is y=-4x+c because it passes through (-8,9) 9=-4*-8+c c=9-32=-23 reqd. eq. of line is y=-4x-23